# Classification of left invariant Riemannian metrics on nonunimodular 4-dimensional Lie groups

### v2

@@ 1.  Introduction
Where $A^{\ast}\langle.,.\rangle$ is the pullback of the inner product .,$\langle$.,$\rangle$ . , . ⟩ under $A$, and hence .,$\langle$.,$\rangle$ . , . ⟩ and $A^{\ast}\langle.,.\rangle$ are isometric. The moduli space of left invariant metrics on a Lie group $G$ with Lie algebra $\mathfrak{g}$ is the orbit space of the action of $\operatorname{Aut}(\mathfrak{g})$ on $\widetilde{\mathfrak{M}}$ given by [11]
@@
In [15], Van Thuong provided a crucial method for classifying inner products on 4-dimensional Lie algebras, focusing on the unimodular case. In this paper, we extend this classification to the nonunimodular case. Below, we provide a comprehensive list of all 4-dimensional nonunimodular Lie algebras [12]
@@
$\operatorname{A}_{2}\oplus\operatorname{2A}_{1}$ $[e_{1},e_{2}]=e_{2}$
$\operatorname{2A}_{2}$ $[e_{1},e_{2}]=e_{2},[e_{3},e_{4}]=e_{4}$
$\operatorname{A}_{3,2}\oplus\operatorname{A}_{1}$ $[e_{1},e_{3}]=e_{1},[e_{2},e_{3}]=e_{1}+e_{2}$
$\operatorname{A}_{3,3}\oplus\operatorname{A}_{1}$ $[e_{1},e_{3}]=e_{1},[e_{2},e_{3}]=e_{2}$
$\operatorname{A}_{3,5}^{\alpha}\oplus\operatorname{A}_{1},\newline 0<\left|\alpha\right|<1$ $[e_{1},e_{3}]=e_{1},[e_{2},e_{3}]=\alpha e_{2}$
$\operatorname{A}_{3,7}^{\alpha}\oplus\operatorname{A}_{1},\alpha>0$ $[e_{1},e_{3}]=\alpha e_{1}-e_{2},[e_{2},e_{3}]=e_{1}+\alpha e_{2}$
$\operatorname{A}_{4,2}^{\alpha},\alpha\neq 0,\newline \alpha\neq-2$ $[e_{1},e_{4}]=\alpha e_{1},[e_{2},e_{4}]=e_{2},[e_{3},e_{4}]=e_{2}+e_{3}$
$\operatorname{A}_{4,3}$ $[e_{1},e_{4}]=e_{1},[e_{3},e_{4}]=e_{2}$
$\operatorname{A}_{4,4}$ $[e_{1},e_{4}]=e_{1},[e_{2},e_{4}]=e_{1}+e_{2},[e_{3},e_{4}]=e_{2}+e_{3}$
@@
$\operatorname{A}_{4,6}^{\alpha,\beta},\alpha\neq 0,\newline \beta\geq 0,\alpha\neq-2\beta$ $[e_{1},e_{4}]=\alpha e_{1},[e_{2},e_{4}]=\beta e_{2}-e_{3},[e_{3},e_{4}]=e_{2}+\beta e_{3}$
$\operatorname{A}_{4,7}$ $[e_{2},e_{3}]=e_{1},[e_{1},e_{4}]=2e_{1},[e_{2},e_{4}]=e_{2},[e_{3},e_{4}]=e_{2}+e_{3}$
$\operatorname{A}_{4,9}^{\beta},-1<\beta\leq 1$ $[e_{2},e_{3}]=e_{1},[e_{1},e_{4}]=(1+\beta)e_{1},[e_{2},e_{4}]=e_{2},\newline [e_{3},e_{4}]=\beta e_{3}$
$\operatorname{A}_{4,11}^{\alpha},\alpha>0$ $[e_{2},e_{3}]=e_{1},[e_{1},e_{4}]=2\alpha e_{1},[e_{2},e_{4}]=\alpha e_{2}-e_{3},\newline [e_{3},e_{4}]=e_{2}+\alpha e_{3}$
$\operatorname{A}_{4,12}$ $[e_{1},e_{3}]=e_{1},[e_{2},e_{3}]=e_{2},[e_{1},e_{4}]=-e_{2},[e_{2},e_{4}]=e_{1}$
@@ 2.  Preliminaries
Let =$\mathcal{B}=\left\{e_{1},...,e_{n}\right\}$ be the canonical basis of $\mathbb{R}^{n}$. The Gram-Schmidt procedure produces an orthonormal basis =$\mathscr{B}=\left\{v_{1},...,v_{n}\right\}$ of the Euclidean space $\left(\mathbb{R}^{n},\langle.,.\rangle\right)$ such that
$[v_{1},...,v_{n}]=X\in Tsup_{n}.$
This means that $Xe_{i}=v_{i}$. Since =$\mathscr{B}=\{v_{1},...,v_{n}\}$ is an orthonormal basis of $(\mathbb{R}^{n},\langle.,.\rangle)$, then
@@
The moduli space of left invariant metrics given by $\mathfrak{M}=\operatorname{Aut}(\mathfrak{g})\backslash\widetilde{\mathfrak{M}}$ is identified with the double coset space $\operatorname{Aut}(\mathfrak{g})\backslash\operatorname{GL}(n,\mathbb{R})/\operatorname{O}(n,\mathbb{R})$. See [15].
@@
    Consider an orthonormal basis for $\operatorname{Aut}(\mathfrak{g})$, defined by an upper triangular matrix with positive diagonal entries :
@@
    Note that $[B^{\prime}],[B^{\prime\prime}]\in\operatorname{GL}(4,\mathbb{R})/\operatorname{O}(4,\mathbb{R})$ lie in the same $\operatorname{Aut}(\mathfrak{g})$ orbit if and only if $AB^{\prime}=B^{\prime\prime}U$, for some $U\in\operatorname{O}(4,\mathbb{R})$, or
@@
Let $M=\operatorname{GL}(2,\mathbb{R})/\operatorname{O}(2,\mathbb{R})$. Consider the subgroup of $\operatorname{GL}(2,\mathbb{R})$ isomorphic to (×$\left(\mathbb{R}^{\times}\right)^{2}\rtimes\mathbb{Z}_{2}$:
@@
Then (×$\left(\mathbb{R}^{\times}\right)^{2}\rtimes\mathbb{Z}_{2}\backslash M\cong\left\{\begin{bmatrix}1&x\\ 0&1\end{bmatrix}|\quad x\geq 0\right\}.$
@@
Let $M=\operatorname{GL}(2,\mathbb{R})/\operatorname{O}(2,\mathbb{R})$. Then identifying +=$\mathbb{R}^{+}=\left\{\begin{bmatrix}a&0\\ 0&a\end{bmatrix}|\quad a\in\mathbb{R}^{+}\right\}.$
+×$\mathbb{R}^{+}\times\operatorname{O}(2,\mathbb{R})$ acts on $M$ by left multiplication, and
(+×$\left(\mathbb{R}^{+}\times\operatorname{O}(2,\mathbb{R})\right)\backslash M\cong\left\{\begin{bmatrix}1&0\\ 0&x\end{bmatrix}|\quad 0<x\leq 1\right\}.$
Where +\$\mathbb{R}^{+}\backslash M$ is identified with the upper half plane.
@@ 3.1 The Lie algebra $\operatorname{A}_{2}\oplus\operatorname{2A}_{1}$
The Lie algebra $\mathfrak{g}=\operatorname{A}_{2}\oplus\operatorname{2A}_{1}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the only nonzero bracket is $[e_{1},e_{2}]=e_{2}$.
@@
$X_{1}=b_{11}e_{1},\quad X_{2}=b_{12}e_{1}+e_{2},\quad X_{3}=b_{13}e_{1}+b_{23}e_{2}+e_{3},\quad X_{4}={\color[rgb]{0,0,1}b_{24}e_{2}}+e_{4}$
@@
$\begin{bmatrix}u\\ v\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}$$\Leftrightarrow$ $a_{6}\begin{bmatrix}b_{23}\\ b_{24}\end{bmatrix}=\begin{bmatrix}a_{11}&a_{15}\\ a_{12}&a_{16}\end{bmatrix}\begin{bmatrix}c_{23}\\ c_{24}\end{bmatrix}$
$\begin{bmatrix}y\\ z\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}$$\Leftrightarrow$ $\hskip 13.6572pt\begin{bmatrix}b_{13}\\ b_{14}\end{bmatrix}=\begin{bmatrix}a_{11}&a_{15}\\ a_{12}&a_{16}\end{bmatrix}\begin{bmatrix}c_{13}\\ c_{14}\end{bmatrix}.$
We can choose an orthogonal matrix $\begin{bmatrix}a_{11}&a_{12}\\ a_{15}&a_{16}\end{bmatrix}$ such that its transpose maps the vector $(c_{13},c_{14})$ to a point on the nonnegative $x$-axis. Therefore, we may assume that $b_{14}=0$ and $b_{13}\geq 0$. However, with that choice, we cannot control the vector $a_{6}(b_{23},b_{24})$. Still, by choice of $a_{6}$, we can ensure that $b_{12}\geq 0$.
@@ 3.2 The Lie algebra $\operatorname{2A}_{2}$
The Lie algebra $\operatorname{2A}_{2}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the only nonzero brackets are
$[e_{1},e_{2}]=e_{2},\qquad[e_{3},e_{4}]=e_{4}.$
@@
The matrix $C^{-1}AB$ is orthogonal precisely when $b_{11}=c_{11}$, $b_{33}=c_{33}$, $a_{6}$ and $a_{16}$ are equal to ±$\pm 1$ and $x=y=z=u=v=0$. This implies that
@@ 3.3 The Lie algebra $\operatorname{A}_{3,2}\oplus\operatorname{A}_{1}$
The Lie algebra $\operatorname{A}_{3,2}\oplus\operatorname{A}_{1}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{1},e_{3}]=e_{1},\quad[e_{2},e_{3}]=e_{1}+e_{2}.$
@@ 3.4 The Lie algebra $\operatorname{A}_{3,3}\oplus\operatorname{A}_{1}$
The Lie algebra $\operatorname{A}_{3,3}\oplus\operatorname{A}_{1}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{1},e_{3}]=e_{1},\quad[e_{2},e_{3}]=e_{2}.$
@@ 3.5 The Lie algebra $\operatorname{A}_{3,5}^{\alpha}\oplus\operatorname{A}_{1}$
The Lie algebra $\operatorname{A}_{3,5}^{\alpha}\oplus\operatorname{A}_{1}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{1},e_{3}]=e_{1},\quad[e_{2},e_{3}]=\alpha e_{2},\quad\text{where}\;0<|\alpha|<1.$
@@
The matrix $C^{-1}AB$ is orthogonal precisely when $a_{1},a_{6}$ and $a_{16}$ are equal to ±$\pm 1$, $b_{33}=c_{33}$ and $x=y=z=u=0$. Thus
@@ 3.6 The Lie algebra $\operatorname{A}_{3,7}^{\alpha}\oplus\operatorname{A}_{1}$
The Lie algebra $\operatorname{A}_{3,7}^{\alpha}\oplus\operatorname{A}_{1}$ has a basis =$\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$[e_{1},e_{3}]=\alpha e_{1}-e_{2},\quad[e_{2},e_{3}]=e_{1}+\alpha e_{2},\quad\text{where}\;\alpha>0.$
@@
If $10<b_{22}<$ 1, then the condition that the upper $2\times 2$ block of $C^{-1}AB$ is orthogonal implies that $a_{2}=0$. Thus $\begin{bmatrix}b_{14}\\ b_{24}\end{bmatrix}=\pm\begin{bmatrix}c_{14}\\ c_{24}\end{bmatrix}.$

### v3

@@ 1.  Introduction
Where $A^{\ast}\langle.,.\rangle$ is the pullback of the inner product $\langle.,.\rangle$ under $A$, and hence $\langle.,.\rangle$ and $A^{\ast}\langle.,.\rangle$ are isometric. The moduli space of left invariant metrics on a Lie group $G$ with Lie algebra $\mathfrak{g}$ is the orbit space of the action of $\operatorname{Aut}(\mathfrak{g})$ on $\widetilde{\mathfrak{M}}$ given by [11]
@@
In [16], Van Thuong provided a crucial method for classifying inner products on 4-dimensional Lie algebras, focusing on the unimodular case. In this paper, we extend this classification to the nonunimodular case. Below, we provide a comprehensive list of all 4-dimensional nonunimodular Lie algebras [13, 12]
@@
$\operatorname{A}_{2}\oplus\operatorname{2A}_{1}$ $e_{1},e_{2}]=e_{2}$
$\operatorname{2A}_{2}$ $e_{1},e_{2}]=e_{2},[e_{3},e_{4}]=e_{4}$
$\operatorname{A}_{3,2}\oplus\operatorname{A}_{1}$ $e_{1},e_{3}]=e_{1},[e_{2},e_{3}]=e_{1}+e_{2}$
$\operatorname{A}_{3,3}\oplus\operatorname{A}_{1}$ $e_{1},e_{3}]=e_{1},[e_{2},e_{3}]=e_{2}$
$\operatorname{A}_{3,5}^{\alpha}\oplus\operatorname{A}_{1},\newline 0<\left|\alpha\right|<1$ $e_{1},e_{3}]=e_{1},[e_{2},e_{3}]=\alpha e_{2}$
$\operatorname{A}_{3,7}^{\alpha}\oplus\operatorname{A}_{1},\alpha>0$ $e_{1},e_{3}]=\alpha e_{1}-e_{2},[e_{2},e_{3}]=e_{1}+\alpha e_{2}$
$\operatorname{A}_{4,2}^{\alpha},\alpha\neq 0,\newline \alpha\neq-2$ $e_{1},e_{4}]=\alpha e_{1},[e_{2},e_{4}]=e_{2},[e_{3},e_{4}]=e_{2}+e_{3}$
$\operatorname{A}_{4,3}$ $e_{1},e_{4}]=e_{1},[e_{3},e_{4}]=e_{2}$
$\operatorname{A}_{4,4}$ $e_{1},e_{4}]=e_{1},[e_{2},e_{4}]=e_{1}+e_{2},[e_{3},e_{4}]=e_{2}+e_{3}$
@@
$\operatorname{A}_{4,6}^{\alpha,\beta},\alpha\neq 0,\newline \beta\geq 0,\alpha\neq-2\beta$ $e_{1},e_{4}]=\alpha e_{1},[e_{2},e_{4}]=\beta e_{2}-e_{3},[e_{3},e_{4}]=e_{2}+\beta e_{3}$
$\operatorname{A}_{4,7}$ $e_{2},e_{3}]=e_{1},[e_{1},e_{4}]=2e_{1},[e_{2},e_{4}]=e_{2},[e_{3},e_{4}]=e_{2}+e_{3}$
$\operatorname{A}_{4,9}^{\beta},-1<\beta\leq 1$ $e_{2},e_{3}]=e_{1},[e_{1},e_{4}]=(1+\beta)e_{1},[e_{2},e_{4}]=e_{2},\newline [e_{3},e_{4}]=\beta e_{3}$
$\operatorname{A}_{4,11}^{\alpha},\alpha>0$ $e_{2},e_{3}]=e_{1},[e_{1},e_{4}]=2\alpha e_{1},[e_{2},e_{4}]=\alpha e_{2}-e_{3},\newline [e_{3},e_{4}]=e_{2}+\alpha e_{3}$
$\operatorname{A}_{4,12}$ $e_{1},e_{3}]=e_{1},[e_{2},e_{3}]=e_{2},[e_{1},e_{4}]=-e_{2},[e_{2},e_{4}]=e_{1}$
@@ 2.  Preliminaries
Let $\mathcal{B}=\left\{e_{1},...,e_{n}\right\}$ be the canonical basis of $\mathbb{R}^{n}$. The Gram-Schmidt procedure produces an orthonormal basis $\mathscr{B}=\left\{v_{1},...,v_{n}\right\}$ of the Euclidean space $\left(\mathbb{R}^{n},\langle.,.\rangle\right)$ such that
$v_{1},...,v_{n}]=X\in Tsup_{n}.$
This means that $Xe_{i}=v_{i}$. Since $\mathscr{B}=\{v_{1},...,v_{n}\}$ is an orthonormal basis of $(\mathbb{R}^{n},\langle.,.\rangle)$, then
@@
[…]
@@
    Consider an orthonormal basis for $\mathfrak{g}$, defined by an upper triangular matrix with positive diagonal entries :
@@
    Note that $B^{\prime}],[B^{\prime\prime}]\in\operatorname{GL}(4,\mathbb{R})/\operatorname{O}(4,\mathbb{R})$ lie in the same $\operatorname{Aut}(\mathfrak{g})$ orbit if and only if $AB^{\prime}=B^{\prime\prime}U$, for some $U\in\operatorname{O}(4,\mathbb{R})$, or
@@
Let $M=\operatorname{GL}(2,\mathbb{R})/\operatorname{O}(2,\mathbb{R})$. Consider the subgroup of $\operatorname{GL}(2,\mathbb{R})$ isomorphic to $\left(\mathbb{R}^{\times}\right)^{2}\rtimes\mathbb{Z}_{2}$:
@@
Then $\left(\mathbb{R}^{\times}\right)^{2}\rtimes\mathbb{Z}_{2}\backslash M\cong\left\{\begin{bmatrix}1&x\\ 0&1\end{bmatrix}|\quad x\geq 0\right\}.$
@@
Let $M=\operatorname{GL}(2,\mathbb{R})/\operatorname{O}(2,\mathbb{R})$. Then identifying $\mathbb{R}^{+}=\left\{\begin{bmatrix}a&0\\ 0&a\end{bmatrix}|\quad a\in\mathbb{R}^{+}\right\}.$
$\mathbb{R}^{+}\times\operatorname{O}(2,\mathbb{R})$ acts on $M$ by left multiplication, and
$\left(\mathbb{R}^{+}\times\operatorname{O}(2,\mathbb{R})\right)\backslash M\cong\left\{\begin{bmatrix}1&0\\ 0&x\end{bmatrix}|\quad 0<x\leq 1\right\}.$
[…]
@@ 3.1 The Lie algebra $\operatorname{A}_{2}\oplus\operatorname{2A}_{1}$
The Lie algebra $\mathfrak{g}=\operatorname{A}_{2}\oplus\operatorname{2A}_{1}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the only nonzero bracket is $e_{1},e_{2}]=e_{2}$.
@@
$X_{1}=b_{11}e_{1},\quad X_{2}=b_{12}e_{1}+e_{2},\quad X_{3}=b_{13}e_{1}+b_{23}e_{2}+e_{3},\quad X_{4}=b_{24}e_{2}+e_{4}$
@@
$\begin{bmatrix}u\\ v\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix} \displaystyle\Leftrightarrow$ $a_{6}\begin{bmatrix}b_{23}\\ b_{24}\end{bmatrix}=\begin{bmatrix}a_{11}&a_{15}\\ a_{12}&a_{16}\end{bmatrix}\begin{bmatrix}c_{23}\\ c_{24}\end{bmatrix}$
$\begin{bmatrix}y\\ z\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix} \displaystyle\Leftrightarrow$ $\hskip 13.6572pt\begin{bmatrix}b_{13}\\ b_{14}\end{bmatrix}=\begin{bmatrix}a_{11}&a_{15}\\ a_{12}&a_{16}\end{bmatrix}\begin{bmatrix}c_{13}\\ c_{14}\end{bmatrix}$.
We can choose an orthogonal matrix $\begin{bmatrix}a_{11}&a_{12}\\ a_{15}&a_{16}\end{bmatrix}$ such that its transpose maps the vector $(c_{13},c_{14})$ to a point on the nonnegative $x$-axis. Therefore, we may assume that $b_{14}=0$ and $b_{13}\geq 0$. However, with that choice, we cannot control the vector $a_{6}(b_{23},b_{24})$. Still, by choice of $a_{6}$, we can ensure that $b_{12}\geq 0$. ∎
Remark 3.2.
If we consider the Riemannian metrics as lower triangular matrices with positive diagonal entries, then we can reduce the dimension of the moduli space of the Lie algebra $\operatorname{A}_{2}\oplus\operatorname{2A}_{1}$ to $\operatorname{dim}(\mathfrak{M})=3$. Because we can find an automorphism of $\operatorname{A}_{2}\oplus\operatorname{2A}_{1}$ such that
$\begin{bmatrix}1&0&0&0\\ a_{5}&a_{6}&0&0\\ a_{9}&0&a_{11}&a_{12}\\ a_{13}&0&a_{15}&a_{16}\end{bmatrix}\begin{bmatrix}b_{11}&0&0&0\\ b_{21}&b_{22}&0&0\\ b_{31}&b_{32}&b_{33}&0\\ b_{41}&b_{42}&b_{43}&b_{44}\end{bmatrix}=\begin{bmatrix}\ast&0&0&0\\ 0&1&0&0\\ 0&\ast&1&0\\ 0&\ast&0&1\end{bmatrix}$.
@@ 3.2 The Lie algebra $\operatorname{2A}_{2}$
The Lie algebra $\operatorname{2A}_{2}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the only nonzero brackets are
$e_{1},e_{2}]=e_{2},\qquad[e_{3},e_{4}]=e_{4}.$
@@
The matrix $C^{-1}AB$ is orthogonal precisely when $b_{11}=c_{11}$, $b_{33}=c_{33}$, $a_{6}$ and $a_{16}$ are equal to $\pm 1$ and $x=y=z=u=v=0$. This implies that
@@ 3.3 The Lie algebra $\operatorname{A}_{3,2}\oplus\operatorname{A}_{1}$
The Lie algebra $\operatorname{A}_{3,2}\oplus\operatorname{A}_{1}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{1},e_{3}]=e_{1},\quad[e_{2},e_{3}]=e_{1}+e_{2}.$
@@ 3.4 The Lie algebra $\operatorname{A}_{3,3}\oplus\operatorname{A}_{1}$
The Lie algebra $\operatorname{A}_{3,3}\oplus\operatorname{A}_{1}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{1},e_{3}]=e_{1},\quad[e_{2},e_{3}]=e_{2}.$
@@ 3.5 The Lie algebra $\operatorname{A}_{3,5}^{\alpha}\oplus\operatorname{A}_{1}$
The Lie algebra $\operatorname{A}_{3,5}^{\alpha}\oplus\operatorname{A}_{1}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{1},e_{3}]=e_{1},\quad[e_{2},e_{3}]=\alpha e_{2},\quad\text{where}\;0<|\alpha|<1.$
@@
The matrix $C^{-1}AB$ is orthogonal precisely when $a_{1},a_{6}$ and $a_{16}$ are equal to $\pm 1$, $b_{33}=c_{33}$ and $x=y=z=u=0$. Thus
@@ 3.6 The Lie algebra $\operatorname{A}_{3,7}^{\alpha}\oplus\operatorname{A}_{1}$
The Lie algebra $\operatorname{A}_{3,7}^{\alpha}\oplus\operatorname{A}_{1}$ has a basis $\mathcal{B}=\left\{e_{1},e_{2},e_{3},e_{4}\right\}$ such that the nonzero brackets are
$e_{1},e_{3}]=\alpha e_{1}-e_{2},\quad[e_{2},e_{3}]=e_{1}+\alpha e_{2},\quad\text{where}\;\alpha>0.$
@@
If $0<b_{22}<1$, then the condition that the upper $2\times 2$ block of $C^{-1}AB$ is orthogonal implies that $a_{2}=0$. Thus $\begin{bmatrix}b_{14}\\ b_{24}\end{bmatrix}=\pm\begin{bmatrix}c_{14}\\ c_{24}\end{bmatrix}$.